A bucket is in the shape of a truncated cone.The height of the bucket is 45 cm,and the diameters of the top and the bottom surfaces are 56 cm and 14 cm respectively.Find the the quantity (in litres) of the water the bucket can hold.
Answer by Boreal(15235)

V=(1/3)(r1^2+r1r2+r2^2)h*pi
=(1/3)*45*(28^2+196+49)*pi
=15(1029)*pi
=48490.5 cm^3
This is 48.4905 liters (sorry to forget PI!!)
this was unrounded for pi; 3.14 will produce a slightly different answer